A rectangular water tank measuring 80cm *60cm*60cm is filled from a pipe of Cross section area 1.5 cm^2, the water emerging at 3.2 m/s. HOW long does it take to fill tank?

Grade:9

1 Answers

Nishant Vora IIT Patna
7 years ago
water emerging at 3.2 m/s = 320 cm/s

hence equating the volume flow rate

320*1.5*t = 80*60*60

t = 600 sec = 10 mins